# astroCAMP: A Community Benchmark and Co-Design Framework for Sustainable SKA-Scale Radio Imaging

### v1

@@ Abstract.
The Square Kilometre Array (SKA) project will operate one of the world’s largest continuous scientific data systems, sustaining petascale imaging under strict power caps. Yet current radio-interferometric pipelines utilise only a small fraction of hardware peak performance—typically 4–14%—due to memory and I/O bottlenecks, resulting in poor energy efficiency and high operational and carbon costs. Progress is further limited by the absence of standardised metrics and fidelity tolerances, preventing principled hardware–software co-design and rigorous exploration of quality–efficiency trade-offs. We introduce astroCAMP, a framework for guiding the co-design of next-generation imaging pipelines and sustainable HPC architectures that maximise scientific return within SKA’s operational and environmental limits. astroCAMP provides: (1) a unified, extensible metric suite covering scientific fidelity, computational performance, sustainability, and lifecycle economics; (2) standardised SKA-representative datasets and reference outputs enabling reproducible benchmarking across CPUs, GPUs, and emerging accelerators; and (3) a multi-objective co-design formulation linking scientific-quality constraints to time-, energy-, carbon-to-solution, and total cost of ownership. We release datasets, benchmarking results, and a reproducibility kit, and evaluate co-design metrics for WSClean and IDG on an AMD EPYC 9334 processor and an NVIDIA H100 GPU. Further, we illustrate the use of astroCAMP for heterogeneous CPU–FPGA design-space exploration, and its potential to facilitate the identification of Pareto-optimal operating points for SKA-scale imaging deployments. Last, we make a call to the SKA community to define quantifiable fidelity metrics and thresholds to accelerate principled optimisation for SKA-scale imaging.
@@
copyright: noneconference: arxiv; 15 December 2025;
@@ 1. Introduction
Modern radio interferometers are entering a regime where computing, not photon collection, limits scientific capability. Arrays such as LOFAR, MWA, ASKAP, and the VLA already generate visibility, calibration, and imaging workloads that strain current HPC facilities, and these pressures intensify for the Square Kilometre Array (SKA), the most data-intensive radio observatory ever built. SKA-Low and SKA-Mid—illustrated in Fig. 1—will stream $\sim$8–20 Tb/s of correlated visibilities to two Science Data Processors (SDPs), each expected to operate within a site-level power cap of 5 MW for several decades (dewdney2009ska; broekema2015ska). Because both SDPs are located on carbon-intensive grids in South Africa and Australia, energy efficiency directly determines operational cost, scientific throughput, and long-term sustainability.
@@
Efficiency requirements for SKA SDPs. Public SKA design documents (dewdney2009ska; broekema2015ska) indicate that the Science Data Processor (SDP) must sustain approximately 16–42 PFLOP/s for SKA-Low and 20–72 PFLOP/s for SKA-Mid, depending on observing mode. For a sustained compute requirement $C_{\rm sus}$ and power envelope $P_{\rm cap}$, the facility-level efficiency is $\eta_{\rm facility}=\frac{C_{\rm sus}}{P_{\rm cap}}.$ Only a fraction of this power is available to compute nodes, as modern HPC systems incur 30–40% overheads for networking, memory, storage, and cooling (PUE 1.3–1.4). Adopting $f_{\rm overhead}\in[0.3,0.4]$, the compute-node requirement becomes $\eta_{\rm compute}=\frac{C_{\rm sus}}{P_{\rm cap}(1-f_{\rm overhead})}.$
Fig. 2 summarizes the required efficiency ranges for SKA-Low and SKA-Mid under 1–5 MW power caps. A 2 MW allocation for SKA-Low and SKA-Mid requires tens of GFLOP/s/W, comparable to the average Green500 systems in November 2025. These typically deliver 30–40 GFLOP/s/W. KAIROS, the current Green500 leader, reaches 73.28 GFLOP/s/W (green5002025nov). This is comparable to the most computationally demanding operation case for SKA-Low, under a power cap of 1 MW. However, Green500 benchmarking numbers reflect idealised, compute-bound kernels (LINPACK), whereas SKA imaging workloads are overwhelmingly memory- and I/O-limited (gridding, FFTs, deconvolution). Even with major algorithmic advances—such as $w$-projection (cornwell2008noncoplanar), $w$-stacking (pratley2019fast), faceting, and IDG (van2018image)—radio imaging remains dominated by memory bandwidth, data movement, and irregular access patterns.
The utilisation gap for imaging software deployments. On modern CPU and GPUs, imaging pipelines typically sustain only 4–14 % of peak floating-point performance (veenboer2020radio), revealing a fundamental mismatch between imaging workloads and commodity architectures. This utilization inefficiency can be appreciated by comparing the required SKA-Low and Mid performance with the pink band in Fig. 2. Earlier modelling (broekema2015ska) reported only 10–20% efficiency for FFT- and gridding-dominated workloads, indicating that naïvely scaling existing pipelines to SKA data rates would demand tens of megawatts—far beyond the SDP site power caps. This gap has direct implications for both the feasibility and long-term sustainability of the SKA. Without substantial improvements in software efficiency and hardware utilisation, meeting SKA’s scientific throughput requirements would necessitate excessive overprovisioning of compute infrastructure and inevitably exceed the power budget. The underlying issue is that bandwidth-bound kernels draw significant power even when compute units remain under-utilised, leading to high operational and carbon costs for comparatively little scientific return.
@@
Enabling principled co-design further requires (1) standardised reference datasets; (2) cross-layer metrics integrating performance, energy, carbon, and cost; (3) science-driven image-quality tolerances; and (4) a reproducible framework for evaluating algorithms and systems under realistic SKA constraints. The community lacks such an infrastructure, obscuring the central question for SKA-scale computing: how much scientific fidelity can be traded for performance, energy, and carbon savings without compromising discovery? Other fields have solved similar challenges. Machine learning, for example, overcame a similar lack of standardisation through community benchmarks such as MLPerf (reddi2020mlperf; farrell2021mlperf), unified datasets, accuracy thresholds, and submission protocols. Radio astronomy lacks an equivalent effort. These gaps hinder agile innovation in the domain and prevent researchers and developers from comparing Fig.s of merit without re-running large, expensive experiments.
@@
    Standardised SKA benchmark suite and datasets. We provide visibility datasets, reference dirty images, and parameterised benchmark configurations in the astroCAMP GitHub repository (astroCAMP2025). With this, we aim to enable reproducible, cross-platform evaluation of SKA imaging pipelines and systematic exploration of hardware–algorithm co-design trade-offs.
@@
    A multi-objective co-design formulation. We formalise imaging as an optimisation problem over algorithmic and architectural parameters, with objectives such as time-to-solution, energy-to-solution, and carbon-to-solution under explicit quality, cost, and power constraints.
@@
    A reproducible design-space exploration workflow. Using the PREESM framework (PREESM2017), we demonstrate for a subset of astroCAMP’s metrics suite how astroCAMP can enable multi-objective design-space exploration and the computation of Pareto fronts in the metrics domain space.
Together, the four contributions form the foundation of the first end-to-end methodology for rigorously evaluating and co-designing SKA-scale imaging pipelines with carbon and cost efficiency as first-class objectives.
@@ 2. SKA SDP Sustainability Challenges and Gaps
Why does optimizing efficiency matter? Recent studies highlight the ecological impact of large-scale scientific computing in astronomy (portegies2020ecological; Aujoux_2021; dos2024assessment; knodlseder2022estimate). In response, the SKA Observatory has embedded sustainability and net-zero objectives into its 50-year roadmap, aligned with the UN Sustainable Development Goals, and adopted the CO2 Performance Ladder to support its net-zero transition (skaoSustainability2024; skaoEnvFootprint2024).
@@
Because operational emissions scale directly with compute efficiency, the limitations of current platforms become critical. In practice, low arithmetic intensity and utilisation mean:
  • Scalability failures: bandwidth-bound kernels cannot strong-scale across many CPU/GPU nodes, limiting throughput even as hardware is added.
@@
Energy efficiency, utilisation, and parallel scaling are therefore not just performance metrics but climate performance indicators, coupling algorithmic design and hardware co-design to measurable carbon reduction. A factor-of-two improvement in application-level GFLOP/s/W directly halves SDP operational emissions at fixed science throughput. Moreover, utilization improvement reduces total cost of ownership by lowering both CAPEX and OPEX (alvarez2025ceo).
@@ 2.1. Gaps in Current Evaluation Practice
Metric Category WSClean IDG DDFacet FPGA BIPP WS-S. This
@@
A1 Time-to-solution $T_{c}$
A2 Energy-to-solution $E_{c}$ $\sim$
A3 Throughput $\Theta$ (vis/s) $\sim$ $\sim$
A4 Energy efficiency $\eta_{E}$ $\sim$
@@
A5 Utilisation / occupancy $U$ $\sim$
A6 Memory bandwidth $B_{\mathrm{mem}}$
A7 Peak memory usage $M_{\mathrm{peak}}$ $\sim$ $\sim$ $\sim$ $\sim$ $\sim$
@@
C1 Carbon-to-solution $C_{c}$
C2 Carbon efficiency $\eta_{C}$
@@
E3 Cost efficiency
Recent wide-field imaging studies have delivered significant advances in algorithmic sophistication and numerical fidelity, yet their evaluation methodology remains largely performance-centric. As summarised in Table 1, runtime and dirty-image RMS are consistently reported across WSClean (offringa2014wsclean), BIPP (tolley2025bipp), WS-Snapshot (wu2025performance), IDG (van2018image), FPGA-RP (corda2022reduced), and DDFacet (tasse2023ddfacet; monnier2020parallelisation), but key architecture-level metrics—power, energy-to-solution, energy efficiency, roofline characterisation, bandwidth sensitivity, and hardware utilisation— are either missing or only implicitly analysed. Even in DDFacet, where multi-node parallelisation has been demonstrated (monnier2020parallelisation), no measurements of power or energy per job are reported. This limits and slows quantitative comparison across CPUs, GPUs, FPGAs, and emerging accelerators towards optimizing efficiency metrics.
System-level behaviour exhibits similar fragmentation. Pipeline-stage breakdowns and parallel efficiency are available in DDFacet’s distributed implementation (monnier2020parallelisation), and partially in IDG (van2018image) and FPGA-RP (corda2022reduced), but metrics such as device occupancy, data locality, bandwidth limits, and end-to-end dataflow behaviour are rarely treated systematically. This hinders the identification of system bottlenecks that directly influence scalability and energy efficiency in SKA-scale deployments.
None of the surveyed works reports carbon-to-solution, carbon efficiency, or any economic metrics, despite their increasing relevance given SKA’s strict power and environmental constraints. Overall, the literature reveals a fragmented evaluation landscape in which no existing study jointly assesses architecture-level efficiency, system-level behaviour, algorithmic fidelity, and sustainability. To date, only one imaging study has reported total energy-to-solution (corda2022reduced), but only at the algorithmic kernel level (IDG), not for the entire imaging pipeline (WSClean). astroCAMP addresses this gap by providing a unified, reproducible, and cross-layer metric framework that enables principled algorithm–hardware co-design and supports energy-, carbon-, and fidelity-aware optimisation for large-scale imaging systems.
@@ 3.1. Benchmarking and Co-Design Methodology
1. Profiling & Measurements. Each benchmark is executed under controlled and repeatable conditions. We stabilise system behaviour (e.g., process pinning, NUMA placement, fixed-frequency modes when available) and begin runs only after the node reaches a steady thermal and power state. Experiments are repeated until runtime variance falls below a documented threshold. Rack-level power distribution units (PDUs) provide energy measurements to calculate energy-to-solution $E_{c}$, while platform telemetry (e.g., PMT (corda2022pmt) for CPUs, NVML for GPUs) is timestamp-aligned with PDU traces to validate device-level behaviour. From these measurements we derive time-to-solution $T_{c}$, throughput, and carbon-to-solution $C_{c}$ using region-specific carbon intensity. All scripts, configurations, and reproducibility guidelines are openly available (astroCAMP2025).
2. Co-Design Metrics Evaluation. For each benchmark configuration—including algorithmic parameters (e.g., IDG kernel sizes) and platform choice (CPU, GPU, FPGA)—we assemble a consistent vector of cross-layer metrics defined in Section 3.2. These metrics capture performance, energy, cost, sustainability, and system constraints (e.g., SKA SDP node power caps). This stage performs metric computation only: it normalises and validates measurements across systems, preparing a unified metric vector that serves as input to optimisation. In the current release, system and platform metrics are evaluated using PREESM (PREESM2017), and sustainability/economic metrics via CEO-DC (alvarez2025ceo). Algorithm-level quality metrics are intentionally excluded in this release and form an open call for the community to converge on standardised tolerances.
3. Co-Design Loop & Multi-Objective Optimisation. The metric vectors from Stage 2 feed into a structured design-space exploration (DSE) loop following the formulation in Section 4. The objective is to systematically evaluate alternative algorithm–architecture mappings and solve the resulting multi-objective optimisation problem to produce reproducible Pareto frontiers and trade-off curves. This stage provides the interpretation layer: identifying efficient regions of the design space, exposing performance–energy–quality bottlenecks, and determining when approximate methods or domain-specific accelerators become necessary to meet SKA-scale efficiency targets.

### v2

@@ Abstract.
The Square Kilometre Array (SKA) will operate one of the world’s largest continuous scientific data systems, sustaining petascale imaging under strict power envelopes. Yet current radio-interferometric pipelines typically achieve only 4–14% of hardware peak because of memory and I/O bottlenecks, resulting in high energy, operational, and carbon costs. Progress is further constrained by the absence of standardised cross-layer metrics and survey-level fidelity tolerances for principled hardware–software co-design. We present astroCAMP, a reproducible benchmarking and co-design framework for SKA-scale imaging. astroCAMP contributes: (1) a unified metric suite spanning performance, utilisation, memory/data-movement behavior, sustainability, economics, and scientific fidelity; (2) standardised SKA-representative datasets, reference outputs, and benchmark configurations for reproducible cross-platform evaluation; (3) a multi-objective co-design formulation linking quality constraints to time-, energy-, carbon-, and cost-to-solution; and (4) a reproducible design-space exploration workflow to derive Pareto-optimal operating regions. We release datasets, scripts, benchmark results, and a reproducibility kit, and evaluate WSClean+IDG on an AMD EPYC 9334 CPU and an NVIDIA H100 GPU. The evaluation shows substantial end-to-end orchestration and synchronization bottlenecks despite efficient kernels in active phases, limited CPU strong scaling, and location-dependent carbon/cost efficiency under realistic grid and electricity-price assumptions. We further illustrate the use of astroCAMP for heterogeneous CPU–FPGA design-space exploration, and its potential to facilitate the identification of Pareto-optimal operating points for SKA-scale imaging deployments. Lastly, we call on the SKA community to define quantifiable fidelity metrics and thresholds to accelerate principled optimisation for SKA-scale imaging.
@@
copyright: noneconference: Conference; Switzerland;
@@ 1. Introduction
Modern radio interferometers are entering a regime where computing, not photon collection, limits scientific capability. Arrays such as LOFAR, MWA, ASKAP, and the VLA already generate visibility, calibration, and imaging workloads that strain current HPC facilities, and these pressures intensify for the Square Kilometre Array (SKA), the most data-intensive radio observatory ever built. SKA-Low and SKA-Mid—illustrated in Fig. 1—will stream $\sim$8–20 Tb/s of correlated visibilities to two Science Data Processors (SDPs), each expected to operate within a site-level power cap of 5 MW for several decades (Dewdney et al., 2009; Broekema et al., 2015). Because both SDPs are located on carbon-intensive grids in South Africa and Australia, energy efficiency directly determines operational cost, scientific throughput, and long-term sustainability.
@@
Efficiency requirements for SKA SDPs. Public SKA design documents (Dewdney et al., 2009; Broekema et al., 2015) indicate that the Science Data Processor (SDP) must sustain approximately 16–42 PFLOP/s for SKA-Low and 20–72 PFLOP/s for SKA-Mid, depending on observing mode [^1] Online SKA SDP performance-model references: https://developer.skao.int/projects/sdp-par-model/en/latest/ and https://ska-telescope.gitlab.io/sdp/ska-sdp-par-model/notebooks/SKA1_Imaging_Performance_Model.html.. For a sustained compute requirement $C_{\rm sus}$ and power envelope $P_{\rm cap}$, the facility-level efficiency is $\eta_{\rm facility}=\frac{C_{\rm sus}}{P_{\rm cap}}.$ Only a fraction of this power is available to compute nodes, as modern HPC systems incur 30–40% overheads for networking, memory, storage, and cooling (PUE 1.3–1.4). Adopting $f_{\rm overhead}\in[0.3,0.4]$, the compute-node requirement becomes $\eta_{\rm compute}=\frac{C_{\rm sus}}{P_{\rm cap}(1-f_{\rm overhead})}.$
Fig. 2 summarizes the required efficiency ranges for SKA-Low and SKA-Mid under 1–5 MW power caps. A 2 MW allocation for SKA-Low and SKA-Mid requires tens of GFLOP/s/W, comparable to the average Green500 systems in November 2025. These typically deliver 30–40 GFLOP/s/W. KAIROS, the current Green500 leader, reaches 73.28 GFLOP/s/W (TOP500.org, 2025). This is comparable to the most computationally demanding operation case for SKA-Low, under a power cap of 1 MW. However, Green500 benchmarking numbers reflect idealised, compute-bound kernels (LINPACK), whereas SKA imaging workloads (non uniform 2D and 3D FFTs) are in theory memory-limited and in practical implementations I/O-limited. Prior works in major algorithmic advances—such as $w$-projection (Cornwell et al., 2008), $w$-stacking (Pratley et al., 2019), faceting, and IDG (Van der Tol et al., 2018b)—radio imaging remain dominated by memory bandwidth (FFTs), data movement, and irregular access patterns.
The utilisation gap for imaging software deployments. On modern CPU and GPUs, imaging pipelines typically sustain only 4–14 % of peak floating-point performance (Veenboer and Romein, 2020), revealing a fundamental mismatch between imaging workloads and commodity architectures, as also indicated by the pink band in Fig. 2. Earlier modelling (Broekema et al., 2015) reported only 10–20% efficiency for FFT- and gridding-dominated workloads, implying that naïvely scaling current pipelines to SKA data rates would require tens of megawatts, well beyond SDP site caps. Without major improvements in software efficiency and hardware utilisation, meeting SKA throughput would require substantial overprovisioning and likely exceed the power budget. The core issue is that inefficient implementations can draw significant power even when compute units are under-utilised, producing high operational and carbon costs for limited scientific return.
@@
Principled co-design further requires standardised datasets, cross-layer metrics, science-driven image-quality tolerances, and reproducible evaluation methods under representative SKA workloads. Without such framework, optimisation risks becoming a blind pursuit of efficiency gains, obscuring the central question for SKA-scale computing: how much scientific fidelity can be traded for performance, energy, and carbon savings without compromising discovery?
Machine learning, for example, overcame a similar lack of standardisation through community benchmarks such as MLPerf (Reddi et al., 2020; Farrell et al., 2021), unified datasets, accuracy thresholds, and submission protocols. In large-scale computational physics, mixed-precision solvers exploit application-level convergence tolerances to improve performance while preserving required scientific accuracy (Clark et al., 2010). In climate modeling, the Coupled Model Intercomparison Project (CMIP) defines standardized experimental protocols and evaluation criteria to ensure comparability across models (Eyring et al., 2016). Radio astronomy lacks an equivalent effort. These gaps hinder agile innovation in the domain and prevent researchers and developers from comparing figures of merit without re-running large, expensive experiments.
@@
    Standardised SKA benchmark suite and datasets. We provide visibility datasets, reference dirty images, and parameterised benchmark configurations in the astroCAMP GitHub repository (SEAMS Project, 2025). With this, we aim to enable reproducible, cross-platform evaluation of SKA imaging pipelines and systematic exploration of hardware–algorithm co-design trade-offs.
@@
    A multi-objective co-design formulation. We formalise imaging as an optimisation problem over algorithmic and architectural parameters, with the objective to minimize time- energy- and carbon-to-solution under explicit quality, cost, and power constraints.
@@
    A reproducible design-space exploration workflow. Using the PREESM framework (Suriano et al., 2017), we demonstrate for a subset of astroCAMP’s metrics suite how astroCAMP can enable multi-objective design-space exploration and the computation of Pareto fronts in the metrics domain space.
Together, the four contributions form the foundation of the first end-to-end methodology for rigorously evaluating and co-designing SKA-scale imaging pipelines with carbon efficiency as first-class objective.
@@ 2. SKA SDP Sustainability Challenges and Gaps
Why does optimizing efficiency matter? Recent studies highlight the ecological impact of large-scale scientific computing in astronomy (Portegies Zwart, 2020; Aujoux et al., 2021; dos Santos Ilha et al., 2024; Knödlseder et al., 2022). In response, the SKA Observatory has embedded sustainability and net-zero objectives into its 50-year roadmap, aligned with the UN Sustainable Development Goals, and adopted the CO2 Performance Ladder to support its net-zero transition (SKAO Communications, 2024b, a).
@@
Because operational emissions scale directly with compute efficiency, the limitations of current platforms become critical. In practice, for SKA-scale workloads, low arithmetic intensity and utilisation mean:
  • Scalability failures: many imaging workloads encounter strong-scaling limits in practice due to communication, synchronization, and serial overheads, yielding diminishing throughput gains as hardware is added.
@@
Energy efficiency, utilisation, and parallel scaling are therefore not just performance metrics but climate performance indicators, coupling algorithmic design and hardware co-design to measurable carbon reduction. A factor-of-two improvement in application-level GFLOP/s/W directly halves SDP operational emissions at fixed science throughput. Moreover, utilization improvement reduces total cost of ownership by lowering both CAPEX and OPEX (Rodríguez Álvarez et al., 2025).
@@ 2.1. Gaps in Current Evaluation Practice
Metric Category WSClean IDG DDFacet FPGA BIPP WS-S. A.CAMP
@@
Time-to-solution $T_{c}$
Energy-to-solution $E_{c}$ $\sim$
Throughput $\Theta$ (vis/s) $\sim$ $\sim$
Energy efficiency $\eta_{E}$ $\sim$
@@
Utilisation / occupancy $U$ $\sim$
Memory bandwidth $B_{\mathrm{mem}}$
Memory efficiency $\eta_{mem}$ $\sim$
Peak memory usage $M_{\mathrm{peak}}$ $\sim$ $\sim$ $\sim$ $\sim$ $\sim$
@@
Carbon-to-solution $C_{c}$
Carbon efficiency $\eta_{C}$
@@
Cost per job
Recent wide-field imaging studies have delivered significant advances in algorithmic sophistication and numerical fidelity, yet their evaluation methodology remains largely performance-centric. As summarised in Table 1, runtime and dirty-image RMS are consistently reported across WSClean (Offringa et al., 2014), BIPP (Tolley et al., 2025), WS-Snapshot (Wu et al., 2025), IDG (Van der Tol et al., 2018b), FPGA-RP (Corda et al., 2022b), and DDFacet (Tasse et al., 2023; Monnier et al., 2020), but key architecture-level metrics—power, energy-to-solution, energy efficiency, roofline characterisation, bandwidth sensitivity, and hardware utilisation— are either missing or only implicitly analysed. Even in DDFacet, where multi-node parallelisation has been demonstrated (Monnier et al., 2020), no measurements of power or energy per job are reported. This limits and slows quantitative comparison across CPUs, GPUs, FPGAs, and emerging accelerators towards optimizing efficiency metrics.
System-level behaviour exhibits similar fragmentation. Pipeline-stage breakdowns and parallel efficiency are available in DDFacet’s distributed implementation (Monnier et al., 2020), and partially in IDG (Van der Tol et al., 2018b) and FPGA-RP (Corda et al., 2022b), but metrics such as device occupancy, data locality, bandwidth limits, and end-to-end dataflow behaviour are rarely treated systematically. This hinders the identification of system bottlenecks that directly influence scalability and energy efficiency in SKA-scale deployments.
None of the surveyed works reports carbon-to-solution, carbon efficiency, or any economic metrics, despite their increasing relevance given SKA’s strict power and environmental constraints. Overall, the literature reveals a fragmented evaluation landscape in which no existing study jointly assesses architecture-level efficiency, system-level behaviour, algorithmic fidelity, and sustainability. To date, only one imaging study has reported total energy-to-solution (Corda et al., 2022b), but only at the algorithmic kernel level (IDG), not for the entire imaging pipeline (WSClean). astroCAMP addresses this gap by providing a unified, reproducible, and cross-layer metric framework that enables principled algorithm–hardware co-design and supports energy-, carbon-, and fidelity-aware optimisation for large-scale imaging systems.
@@ 3.1. Benchmarking and Co-Design Methodology
1. Profiling & Measurements. Each benchmark is executed under controlled and repeatable conditions. We wait for node to reach a steady thermal and power state. Rack-level power distribution units (PDUs) provide energy measurements to calculate energy-to-solution $E_{c}$, while platform telemetry (e.g., PMT (Corda et al., 2022a) for CPUs, NVML for GPUs) is timestamp-aligned with PDU traces to validate device-level behaviour. From these measurements we derive time-to-solution $T_{c}$, throughput, and carbon-to-solution $C_{c}$ using region-specific carbon intensity. All scripts, configurations, and reproducibility guidelines are openly available (SEAMS Project, 2025).
2. Co-Design Metrics Evaluation. For each benchmark configuration—including algorithmic parameters (e.g., IDG kernel sizes) and platform choice (CPU, GPU, FPGA)—we assemble a consistent vector of cross-layer metrics defined in Section 3.2. These metrics capture performance, energy, cost, sustainability, and system constraints (e.g., SKA SDP node power caps). This stage performs metric computation only: it normalises and validates measurements across systems, preparing a unified metric vector that serves as input to optimisation. In the current release, system and platform metrics are evaluated using PREESM (Suriano et al., 2017), and sustainability/economic metrics via CEO-DC (Rodríguez Álvarez et al., 2025). Algorithm-level quality metrics are intentionally excluded in this release and form an open call for the community to converge on standardised tolerances.
3. Co-Design Loop & Multi-Objective Optimisation. The metric vectors from Stage 2 feed into a structured design-space exploration (DSE) loop following the formulation in Section 4. The objective is to systematically evaluate alternative algorithm–architecture mappings and solve the resulting multi-objective optimisation problem to produce reproducible Pareto frontiers and trade-off curves. This stage provides the interpretation layer: identifying efficient regions of the design space, exposing performance–energyquality bottlenecks, and determining when approximate methods or domain-specific accelerators become necessary to meet SKA-scale efficiency targets.